Skip to content
Solved Assignments

Statistical Analysis and Probability Evaluation: Comprehensive Problem Set Solutions

A complete set of solved problems in statistical analysis and probability.

Preview 5 of 16 pages

See the first 5 pages exactly as they appear in the document.

  1. Page 1
    Read page 1 as text

    Statistical Analysis and Probability Evaluation: Comprehensive Problem Set Solutions Name: Lowenbrau Date: 5 /9/2015 Problem Number Solution 1. (a) FALSE (b) FALSE (c) FALSE (d) TRUE (e) FALS E Work: (a) If all the observations in a data set are identical, then its variance is equals to zero. (b) Since A and B are disjoint, the probability of their intersection must be equal to zero, not equal to 0.9 . (c) For a symmetric distribution if P ( X > 2) = 0.3, then P ( X ≤ – 2) = 0.3, not P ( X ≤ – 3) ≤ 0.3 . (d) The critical value for 9 5 9% confidence level is larger than that of 9 0 % confidence level . So 9 5 % confidence interval has larger width than 9 0 % confidence interval. (e) S ince the p - value for this right tailed test, P (T > 1.5) = 1 – P (X < 1.5) = 1 – 0.96 = 0.04 < 0.05, the decision must be reject the null hypothesis is at 0.05 level of significance. .

  2. Page 2
  3. Page 3
    Read page 3 as text

    2. The shaded cells in the following table indicate the calculated values. Check out Time (in minutes) Frequency Relative Frequency 1.0 - 1.9 3 0.12 2.0 - 2.9 12 0.48 3.0 - 3.9 5 0.2 0 4.0 - 4.9 3 0.12 5.0 - 5.9 2 0.08 Total 25 1.00 Work: Total frequency = 25  Value of f when relative frequency is 0.2 0 = 25 X 0.2 0 = 5 . 3 + 12 + 5 + 3 + ? = 25 23 + ? = 25 ? = 25 – 23 = 2 Frequency of the cell Relative frequency Total frequency Relative frequency for 1.0 1.9 3 25 0.12 Relative frequency for 2.0 2.9 12 25 0.48 Relative frequency for 3.0 3.9 0.20 Relative frequency for 4.0               4.9 3 25 0.12 Relative frequency for 5.0 5.9 2 25 0.08       3. Answer: 4 0 % Work: Obtain the percentage of the checkout times was at least 3 minutes, 5 3 2 10 40% 25 25    

  4. Page 4
    Read page 4 as text

    4. Answer: The median must lie in the class 2 .0 – 2 .9. Work: The following table shows the cumulative relative frequency distribution of the given data. Check out Time Frequency Relative Frequency Cumulative Relative Frequency 1.0 - 1.9 3 0.12 0.12 2.0 - 2.9 12 0.48 0.60 3.0 - 3.9 5 0.2 0.80 4.0 - 4.9 3 0.12 0.92 5.0 - 5.9 2 0.08 1.00 Total 25 1.00 - The median of the distribution covers 50% of the observations in the given dataset. Here the cumulative relative frequency corresponding to 2 .0 – 2 .9 is just greater than 50%, so median must lie in that interval. 5. Answer: This distribution has positive skew . Work: Since the mean of the distribution lie in the class 3.0 – 3.9, and it is greater than the median the distribution has positive skew . 6. Answer: The re quired probability = 1/221 . If the card selection is without replacement, the probability that the first card is an ace and the second card is also an ace is, 4 3 1 52 51 221  

  5. Page 5
    Read page 5 as text

    7. Answer: The required probability = 1/169 . Work: If the card selection is with replacement, the probability that the first card is an ace and the second card is also an ace is, 4 4 1 52 52 169   8 . Answer: (c) Quiz 1 has the less value of the interquartile range. Work: T he interquartile range for Quiz 1 = 8 5 – 40 = 4 5 . T he interquartile range for Quiz 2 = 9 0 – 35 = 55 . 9 . Answer: ( a ) Quiz 2 Work: For the Quiz 2 the percentage of students with grades 90 or over is 25%, since Q3 = 90 . For the Quiz 1 the percentage of students with grades 90 or ove r is less than 25%, since Q3 = 8 5, indicates that 25% of students are with grades 8 5 or over . So subtracting the percentage between grades 85 and 9 0 , the required percentage is less than 25%. 1 0 . Answer: ( b ) Quiz 2 Work: For the Quiz 1 the percentage of students with grades 60 or less is more than 50 %, since Q 2 = 55 . For the Quiz 2 the percentage of students with grades 6 0 or less is also mor e than 5 0 %, since Q 2 = 50 . But the percentage between grades 5 0 and 60 , is slightly more than that of the percentage between grades 55 and 60, so Quiz 2 has a greater percentage.

11 more pages in the full document

Unlock it once and it stays in your library, ready to chat with or turn into flashcards.

Unlock the full document
See all

Study this document with CramX

Once it is in your library, every tool can work from it.